For the following data set, calculate the Inter-Quartile Range (IQR).

124 45 52 78 190 120 34 87 111 60 10
Answer #1

Data in ascending order is

10,34,45,52,60,78,87,111,120,124,190

First quartile is median for lower half {10,34,45,52,60}

Q1 = middle term in set ...................because we have odd number of data values

Q1 = 3rd term

Q1 = 45

Similarly,

third quartile is median for lower half {87,111,120,124,190}

Q3 = middle term in set ...................because we have odd number of data values

Q3 = 3rd term

Q3 = 120

Inter quartile range = Q3-Q1

IQR = 120 -45 = 75